OpenStudy (anonymous):

if angle acd=100 and angle b=60 find angle a

OpenStudy (anonymous):



OpenStudy (anonymous):

i thought the same thing @Roh but i only have answer choices 60 40 80 160

OpenStudy (anonymous):

a)60 b)40 c)80 d)160

OpenStudy (anonymous):

i think it will be solved like this : 180-110=70 and then for angle A : 180-(70+60)=50

OpenStudy (anonymous):

@Roh but how come it doesnt show on the awnser choices ??

OpenStudy (anonymous):

are you sure the question is correct??

OpenStudy (anonymous):

yeah thats how the question is written

hero (hero):

B = 60 ACD = 180 ACD + ACB = 180 100 + ACB = 180 ACB = 180 - 100 ACB = 80 B + ACB = 60 + 80 = 140 A = 180 - 140 = 40

OpenStudy (anonymous):

@Hero i get it now thanks !!!!

OpenStudy (jdoe0001):